OpenStudy (anonymous):

What is the value of t in the equation 3(2t + 5) = 5t + 25? 5 6 10 35

OpenStudy (anonymous):

6t+15=5t+25 Do you know how to proceed from here?

OpenStudy (anonymous):

no

OpenStudy (aaronandyson):

" 3(2t + 5) = 5t + 25? " 3(2t + 5) Multiple each term in the bracket by 3 2t*3 and 5*3 Therfore,2t*3 = 6t and 5*3 = 15

OpenStudy (aaronandyson):

Therefore,6t+15=5t+25 Combine Like terms Therefore,6t - 5t = 25 - 15 Therefore, 1t or just t = 10 Therefore, t= 10